# Learning the Exact SABR Model

### v1

@@ Abstract
We develop a novel SABR DNN, a specialized Artificial Deep Neural Network (DNN) architecture that learns the true SABR stochastic dynamics using an unprecedented large training dataset (more than 200 million points) of interest rate Cap/Floor volatility surfaces, including very long maturities (30Y) and extreme strikes consistently with market quotations. Our dataset is obtained via high-precision unbiased Monte Carlo simulation of a special scaled shifted-SABR stochastic dynamics, which allows dimensional reduction without any loss of generality. Our SABR DNN provides arbitrage-free calibration of real market volatility surfaces and Caps/Floors prices for any maturity and strike with negligible computational effort and without retraining across business dates.
Our results fully address the gaps in the previous machine learning SABR literature in a systematic and self-consistent way, and can be extended to cover any interest rate European options in different rate tenors and currencies, thus establishing a comprehensive functional SABR framework that can be adopted for daily trading and risk management activities.
JEL classifications: C45, C63, G12, G13.
Keywords: stochastic volatility; SABR, Hagan, model calibration; volatility surface; interest rate derivatives; Cap; Floor; Monte Carlo simulation; machine learning in finance; deep neural network.
@@ 1.4 Our Contribution
Although the limitations listed above are scattered across different papers, to the best of our knowledge, no single work where all of them are addressed together for the specific SABR model in a systematic and self-consistent way. In this paper, we aim to cover these gaps by combining all the key elements and establish a comprehensive and functional SABR framework that practitioners can adopt in their daily trading and risk management activity on interest rate options. Specifically, we contribute to this research field with the following elements.
@@ 2.2 Scaled Shifted-SABR Model
Consistently, we also rescale the Caplet/Floorlet shifted strike to $\hat{K}=\frac{\bar{K}}{\bar{F}_{0}}$. The proof is provided in app. B.3. Finally, we set [^9] Note that we denote always with $\theta^{\textit{SABR}},\theta^{\textit{CF}}$ different combinations of model and contract parameters, e.g. scaled and not-scaled, hoping that the difference will be clear depending on the context. $\theta^{\textit{SABR}}=\left\{\hat{\alpha},\beta,\rho,\nu\right\}$ and $\theta^{\textit{CF}}=\left\{T,\hat{K}\right\}$. Note that we do not further reduce the model parameters by fixing $\beta=1$ as done in some previous papers, e.g. [31].
@@ 3.1 Datasets
The parameters’ ranges were selected so as to generate a training set spanning a sufficiently wide range of values, enabling an effective and robust calibration of the SABR model in different market conditions, while at the same time avoiding excessively broad ranges that could lead to the inclusion of unrealistic parameter values. The initial forward rate $F_{0}$ was centered around market-observed values, i.e. EURIBOR6M forward rates, and the range was extended to include all levels the rate may reasonably assume under the prevailing economic conditions. The shift parameter was fixed to $\lambda=3\%$, consistently with market quotations for EUR Caps/Floors which include negative strikes (see app. A.2). The parameters $F_{0}$ and $\alpha$ are then rescaled as discussed in sec. 2.2. The ranges for the remaining SABR model parameters reported in tab. 1 were determined using the following empirical iterative procedure.
@@ 3.2 DNNs Setup and Training
Output $\sigma^{\textit{DNN}}_{\textit{SLN}}(\hat{\alpha},\beta,\rho,\nu,T,\hat{K})=\sigma^{\textit{DNN}}_{\textit{SLN}}(\alpha,\beta,\rho,\nu,T,K)$
@@
The input layer has 6 nodes to accommodate the model parameters $\theta^{\textit{SABR}}=\{\hat{\alpha},\beta,\rho,\nu\}$ and the contract parameters $\theta^{\textit{CF}}=\{T,\hat{K}\}$ (the DNNs’ input features). The output layer consists of a single node representing the shifted-Black implied volatility $\sigma^{\textit{DNN}}_{\textit{SLN}}$ with a shift parameter $\lambda=3\%$ consistent with the input data. The linear activation function is used to get the output, since if the neural network is learning properly, it should naturally produce positive implied volatilities, without the need to explicitly enforce it. The DNNs’ input data are standardized using the StandardScaler from scikit-learn, which transforms each input feature to have zero mean and unit variance, to allow faster convergence and greater efficiency. The DNNs’ architecture was optimized empirically, checking that reducing the number of hidden layers or nodes per layer led to a worse performance, while increasing them had little or no effect.
@@ 3.3 Volatility Smile Calibration
Figure 3: Shifted-SABR model calibrations for short (1.5y), medium (10y) and long (30y) maturities. Black crosses: market data. Blue lines: SABR DNN smile (left) obtained from the SABR parameters calibrated with SABR DNN, and SABR Hagan smile (right) obtained from the SABR parameters calibrated with Hagan et al. approximation. Red lines: MC SABR DNN smile (left) obtained by MC simulations using the SABR parameters calibrated with SABR DNN, MC SABR Hagan smile (right) obtained by MC simulations using the SABR parameters calibrated with Hagan et al. approximation. High precision MC simulations performed with $2^{24}$ scenarios and time steps as in tab. 4. MC errors not shown being negligible. The numerical values corresponding to the smiles depicted in the figure and the respective calibrated parameters are provided in app. E.

### v2

@@ Abstract
We develop a novel SABR DNN, a specialized Artificial Deep Neural Network (DNN) architecture that learns the true SABR stochastic dynamics using an unprecedented large training dataset of interest rate Cap/Floor volatility surfaces (more than 200 million points), including very long maturities (30Y) and extreme strikes consistently with market quotations. Our dataset is obtained via high-precision unbiased Monte Carlo simulation of a scaled shifted-SABR stochastic dynamics, which allows dimensional reduction without any loss of generality.
Our SABR DNN provides arbitrage-free calibration of real market volatility surfaces and Cap/Floor prices for any maturity and strike with negligible computational effort and without retraining across business dates. Our results fully address the gaps in the previous machine learning SABR literature in a systematic and self-consistent way, and can be extended to cover any interest rate European options in different rate tenors and currencies, thus establishing a comprehensive functional SABR framework that can be adopted for daily trading and risk management activities.
Classifications: JEL: C45, C63, G12, G13; MSC: 91G60 (Primary) 60H35, 68T07, 65C05 (Secondary); ACM: G.1.6, I.2.6, I.5.1, J.4.
Keywords: stochastic volatility; SABR, Hagan, model calibration; volatility surface; interest rate derivatives; Cap; Floor; Monte Carlo simulation; machine learning; deep neural network.
@@ 1.4 Our Contribution
Although the limitations listed above are scattered across different papers, to the best of our knowledge there is no single work where all of them are addressed together for the specific SABR model in a systematic and self-consistent way. In this paper, we aim to cover these gaps by combining all the key elements and establish a comprehensive and functional SABR framework that practitioners can adopt in their daily trading and risk management activity on interest rate options. Specifically, we contribute to this research field with the following elements.
@@ 2.2 Scaled Shifted-SABR Model
Consistently, we also rescale the Caplet/Floorlet shifted strike to $\hat{K}=\frac{\bar{K}}{\bar{F}_{0}}$. Finally, we set [^9] Note that we denote always with $\theta^{\textit{SABR}},\theta^{\textit{CF}}$ different combinations of model and contract parameters, e.g. scaled and not-scaled, hoping that the difference will be clear depending on the context. $\theta^{\textit{SABR}}=\left\{\hat{\alpha},\beta,\rho,\nu\right\}$ and $\theta^{\textit{CF}}=\left\{T,\hat{K}\right\}$. Note that the Caplet/Floorlet implied volatlity is the same, i.e. $\sigma(\hat{\alpha},\beta,\rho,\nu,T,\hat{K})=\sigma(\bar{F}_{0},\alpha,\beta,\rho,\nu,T,\bar{K})$, The proof is provided in app. B.3.
We stress that we do not further reduce the model parameters by fixing $\beta=1$ as done in some previous papers, e.g. [31].
@@ 3.1 Datasets
The parameters’ ranges were selected so as to generate a training set spanning a sufficiently wide range of values, enabling an effective and robust calibration of the SABR model in different market conditions, while at the same time avoiding excessively broad ranges that could lead to the inclusion of unrealistic parameter values. The initial forward rate $F_{0}$ was centered around market-observed values, i.e. EURIBOR6M forward rates, and the range was extended to include all levels the rate may reasonably assume under the prevailing economic conditions. The shift parameter was fixed to $\lambda=3\%$, consistently with market quotations for EUR Caps/Floors which include negative strikes (see app. A.2). The parameters $F_{0}$ and $\alpha$ are then scaled as discussed in sec. 2.2. The ranges for the remaining SABR model parameters reported in tab. 1 were determined using the following empirical iterative procedure.
@@ 3.2 DNNs Setup and Training
Output $\sigma^{\textit{DNN}}_{\textit{SLN}}(\hat{\alpha},\beta,\rho,\nu,T,\hat{K})=\sigma^{\textit{DNN}}_{\textit{SLN}}(\bar{F}_{0},\alpha,\beta,\rho,\nu,T,\bar{K})$
@@
The input layer has 6 nodes to accommodate the scaled shifted-SABR and contract parameters $\theta=\{\hat{\alpha},\beta,\rho,\nu,T,\hat{K}\}$ (the DNNs’ input features). The output layer consists of a single node representing the shifted-Black implied volatility $\sigma^{\textit{DNN}}_{\textit{SLN}}$ with a shift parameter $\lambda=3\%$ consistent with the input volatility data. Note that the DNN works with scaled shifted-SABR parameters but the output implied volatility is the same as the non-scaled model, as discussed in sec. 2.2. The linear activation function is used to get the output, since if the neural network is learning properly, it should naturally produce positive implied volatilities, without the need to explicitly enforce it. The DNNs’ input data are standardized using the StandardScaler from scikit-learn, which transforms each input feature to have zero mean and unit variance, to allow faster convergence and greater efficiency. The DNNs’ architecture was optimized empirically, checking that reducing the number of hidden layers or nodes per layer led to a worse performance, while increasing them had little or no effect.
@@ 3.3 Volatility Smile Calibration
Figure 3: Shifted-SABR model calibrations for short (1.5y, top panel), medium (10y, mid panel) and long (30y, bottom panel) maturities. Black crosses: market data. Blue lines: SABR DNN smile (left) obtained from the SABR parameters calibrated with SABR DNN, and SABR Hagan smile (right) obtained from the SABR parameters calibrated with Hagan et al. approximation. Red lines: MC SABR DNN smile (left) obtained by MC simulations using the SABR parameters calibrated with SABR DNN, MC SABR Hagan smile (right) obtained by MC simulations using the SABR parameters calibrated with Hagan et al. approximation. High precision MC simulations performed with $2^{24}$ scenarios and time steps as in tab. 4. MC errors not shown being negligible. The same scales are used for all panels to allow direct comparability. The numerical values corresponding to the smiles depicted in the figure and the respective calibrated parameters are provided in app. E.